OpenStudy (sammietaygreen):

Find the values of x for which the equation cos x = -1 is true. k represents an integer.

OpenStudy (anonymous):

there is no k in your question

OpenStudy (anonymous):

7

OpenStudy (sammietaygreen):

I got the last one

OpenStudy (solomonzelman):

cos(x)=-1 cos(x)+1=0 That means that your x-solutions are the roots of y=cos(x)+1 https://www.desmos.com/calculator/ahvced9xlt

OpenStudy (sammietaygreen):

What do you mean by roots

OpenStudy (solomonzelman):

roots of the functions are the zeros (or the x-intercepts) of the function.

OpenStudy (sammietaygreen):

I still don't get it honestly. Do I have to find the zeros of the equation

OpenStudy (sammietaygreen):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

for each answer choice, let's replace k with some whole number k = 0 is probably the easiest to deal with

jimthompson5910 (jim_thompson5910):

so instead of 2pi*k, it is 2pi*k = 2pi*0 = 0 what happens when you plug x = 0 into the original equation?

OpenStudy (sammietaygreen):

well, cos(0) = 1

OpenStudy (sammietaygreen):

is this basically plugging in?

jimthompson5910 (jim_thompson5910):

so `cos(x) = -1` turns into `1 = -1` when you replaced x with 0 and evaluated the left hand side

jimthompson5910 (jim_thompson5910):

`1 = -1` is a false equation, so x = 0 is NOT a solution (it is not in the solution set)

OpenStudy (sammietaygreen):

I plugged in pi + 2pi(k) and got -1

jimthompson5910 (jim_thompson5910):

did you first replace k with 0 and then simplify to get pi?

OpenStudy (sammietaygreen):

No. Oops. So I have to do cos(*insert equation*) and then have k be 0?

jimthompson5910 (jim_thompson5910):

do you understand why I'm replacing k with 0?

OpenStudy (sammietaygreen):

bc it represents an integer?

jimthompson5910 (jim_thompson5910):

yes, k is any integer and 0 is the easiest to work with because 0 makes things go away pi+2pi*k = pi+2pi*0 = pi + 0 = pi so `pi+2pi*k` is equivalent to `pi` when k = 0

jimthompson5910 (jim_thompson5910):

you'll then find `cos(pi) = -1` so... cos(x) = -1 cos(pi) = -1 ... replace x with pi -1 = -1 ... true equation because the last equation is true, the first equation is true when x = pi so x = pi is definitely one of the infinitely many solutions x = pi is in the solution set for the equation cos(x) = -1

OpenStudy (sammietaygreen):

So how do I figure out which one is right?

jimthompson5910 (jim_thompson5910):

do you agree that x = pi is a solution?

OpenStudy (sammietaygreen):

Well, that's one of the solutions I got as well so yes

jimthompson5910 (jim_thompson5910):

ok try k = 1 and plug it into pi+2pi*k what is `pi+2pi*k` equivalent to when k = 1 ?

OpenStudy (sammietaygreen):

9.4 :/

jimthompson5910 (jim_thompson5910):

in terms of pi

OpenStudy (sammietaygreen):

so it would be like 2*1 or 3.14 + 2*1?

jimthompson5910 (jim_thompson5910):

pi+2pi*k = pi + 2pi*1 = ??? treat pi as if it were a variable

OpenStudy (sammietaygreen):

2pi?

jimthompson5910 (jim_thompson5910):

think of it like 2 pies + 1 pie = how many pies?

OpenStudy (sammietaygreen):

3 pies lol

jimthompson5910 (jim_thompson5910):

yep so 2pi+pi = 3pi

jimthompson5910 (jim_thompson5910):

if k = 1, then pi+2pi*k = 3pi

jimthompson5910 (jim_thompson5910):

tell me what cos(x) is equal to when x = 3pi

OpenStudy (sammietaygreen):

-1

jimthompson5910 (jim_thompson5910):

now try k = 2. Plug it into pi+2pi*k to get what?

OpenStudy (sammietaygreen):

So I would just add the pi's the same way?

jimthompson5910 (jim_thompson5910):

yeah it's similar to pi+2pi

OpenStudy (sammietaygreen):

so would it be 4pi?

jimthompson5910 (jim_thompson5910):

pi+2pi*k = pi+2pi*2 = ?????

OpenStudy (sammietaygreen):

6? cause we're keeping it in terms of pi?

OpenStudy (sammietaygreen):

ill brb I have to workout

jimthompson5910 (jim_thompson5910):

pi+2pi*k = pi+2pi*2 = pi+4pi = ?????

OpenStudy (sammietaygreen):

5pi?

OpenStudy (sammietaygreen):

cos(5pi) = -1 too

jimthompson5910 (jim_thompson5910):

yep, let's do one more if k = 3, then pi+2pi*k is equal to what?

OpenStudy (sammietaygreen):

8pi?

jimthompson5910 (jim_thompson5910):

just replace k with 3 and simplify \[\Large \pi+2\pi*\color{red}{k}=\pi+2\pi*\color{red}{3} = ???\]

OpenStudy (sammietaygreen):

wait, no 9

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (sammietaygreen):

wouldn't it be 9pi though? I don't see how it qwouldnt be